OpenStudy (anonymous):

Marna wants to store one hundred fifty cubic inches of homemade pasta sauce in a food storage container. Should she choose the cylinder with a radius of 3 inches and a height of 5 inches, or the cube with a side length of 6 inches? Explain.

OpenStudy (turingtest):

The formula for the volume of a cylinder is\[V=\pi r^2\ell\]where r is the radius and l is the side length which measuerments give the larger volume

OpenStudy (turingtest):

?

OpenStudy (bahrom7893):

pi*r^2*h=pi*9*5=45pi=141.37.. u want 150, so no u won't be able to store

OpenStudy (anonymous):

You would need the 6 in side